Solve using Cramer's Rule
2x-7y=-4
x-y=3

Answer by invaderjmg(2)

for 2 variable system:
A1x + B1y = C
A2x + B2y = C
we define:(use this formulas in 2 var. below for solving the determinants first)
D =
A1 B1
A2 B2

Dx =
C1 B1
C2 B2

Dy =
A1 C1
A2 C2
Then: (using cramers rule for final step)
X = DX/D
Y = DY/D
given(2 Variables):
2x - 7y = -4
x - y = 3
solution: (just follow the formulas above:)

D =
2 -7
1 -1

Dx =
-4 -7
3 -1

Dy =
2 -4
1 3
now calculate each of these determinants:
D = 2(-1) - (-7)(1) = -2 + 7 = 5
Dx = (-4)(-1) - 3(-7) = 4 + 21 = 25
Dy = 3(2) - (-4)1 = 6 + 4 = 10
so now i already find each of their determinants... so we can now use the cramers rule......(its very simple by using the formula above)
X = Dx/D = 25/5 = 5
Y = Dy/D = 10/5 = 2
congratulations...!!! youve now already solved the x & y's !!!! Invader Rulezzz!